# Degree Matrix Comparison for Graph Alignment

### v2

@@ 5.1 Real World Networks
Table 1: Comparison with baselines on PPI network [4, 5, 6, 8]. Deletion probability $p_{d}=0.01$. “Two-Step” uses degrees for alignment, and refines it through average degrees of neighbors. “Euc Dist.” aligns two graphs by minimizing distances between embeddings. DMC has highest accuracy, but larger complexity compared to [5, 6].
(a) Social Networks
(b) Roadmaps
Figure 3: We use the random walk graph sampling method for these experiments.
@@
Table 2: Means and variances of degree distributions of $G_{s}$. The best performing Facebook and YouTube have large variances in their degree distributions. Comparing Flickr, Last.fm, and Myspace, Flickr and Last.fm are much more heterogeneous. However, there is no strict correlation between variance and performance, as can be seen from California and Pennsylvania roadmaps.
Networks Percent Correct Density Avg. Clustering
bio-grid-fission-yeast .9932 .0123 .1874
bn-mouse-retina-1 .9900 .9983 1.5539
bio-ce-gn .9894 .0218 .1839
bn-fly-drosophila-medulla-1 .9378 .0211 3.8729
bio-grid-human .8440 .0014 .0800
bn-mouse-kasthuri .6156 .0032 0
bio-yeast .5822 .0018 .0708
Table 3: Comprehensive table of experiments on biological networks, ordered by DMC’s performance on $p=90\%$ overlap subgraphs. Density highlight threshold: .01, average clustering coefficient highlight threshold: .1. DMC works better on graphs with higher densities and clusterings.
(a) Performance Decay Plot
$\epsilon$ Correct Percentage
0 $96.58\%$
10 $96.32\%$
100 $50.46\%$
1000 $0.18\%$
(b) Greedy DMC Performance
Figure 4: Experiments on PPI, with random deletion sampling method.
5.2 Graphs from Generation Models
5.2.1 Different Graph Generation Models
In this section, we introduce the Erdos-Renyi, Barabasi-Albert, and Chung-Lu models that were derived to generate complex networks, potentially to model real world network behaviors.
The Erdos-Renyi model is the most basic random graph generation model. One common way to define this is to fix the number of vertices, say $n_{er}$, then connect an edge between every pair of vertices with some probability $p_{er}$. The resulting graph is typically denoted as $G(n_{er},p_{er})$ [22]. One interesting property of these graphs is that their degree distributions are binomial, as can be be seen by the following calculation:
@@ 5.2.1 Different Graph Generation Models
The Barabasi-Albert model considers properties observed on real world networks, especially the preferential attachment phenomena [23]. Instead of attaching new edges with uniform probability, it is possible to attach new edges with weighted probabilities. Suppose we already have $m$ nodes with some connections, then for the $(m+1)^{\text{th}}$ node, the probability that it will be attached to a previous node $v_{i}$ is
@@
We can observe that the probabilities of attaching the new node to any previous node sums up to one, forming a probability distribution; moreover, higher degree nodes attract more new nodes. This usually leads to what we call hubs, which are local centers that are attached to many low degree nodes.
Now we introduce the Chung-Lu model. It also attempts to capture real world network properties, but instead of focusing on preferential attachment or community behaviors, they attempt to approximate real world networks’ degree distributions [24]. Suppose we have a specific degree distribution in mind and we expect $m$ nodes in total, then the attachment probability between each node $v_{i}$ and another node $v_{j}$ is as follows:
@@
Moreover, the expected degree at each node $v_{i}$ is
@@
5.2.2 Discussion on Model Approximations to Real World Networks
@@
5.2.3 Applying DMC to Generated Graphs
Definition 5.1 (k-partite Graph).
@@ 5.2.3 Applying DMC to Generated Graphs
We applied DMC to graphs generated through the three models (with some slight variations). We examine the performance of DMC on the graphs with respect to different partite numbers $k$.
@@
For our Erdos-Renyi experiments, we fix total number of nodes to be $n_{er}=100$, then split the nodes into $k$ partite groups; the partition of nodes is arbitrary. After that, we connect edges between different partite groups with probability of $p_{er}=0.5$. Then we use the random walk graph sampling method to obtain two graphs with 90% overlap, and perform DMC. We run the same experiment for 20 times, then take the mean performance.
For our Barabasi-Albert experiments, we do a similar thing. Every time we add a new node, we connect it to five different previous nodes (add 5 edges). For our Chung-Lu experiments, we use a power-law distribution as our initial degree distribution, where the power-law exponent was chosen to be $\gamma=2.5$, and we set a maximum degree limit of 20.
Figure 5: Discovery of negative linear relationship.
Figure 6: BA model has much better performance than ER. (It is more heterogeneous.)
Figure 7: CL model with initial power-law degree distribution.
@@
Real world networks are more complex than we can imagine, hence the difficulty of capturing them with mathematical language. However, it is interesting the simple DMC works better on the more complex real world networks than synthetically generated graphs. This shows that modeling complex networks is no easy task, and understanding the complexity will potentially bring us rich information.
@@ 6 Conclusion and Future Work
We proposed the DMC and Greedy DMC, showed that they were theoretically motivated, and analyzed experimental results. DMC’s accuracy on classic PPI networks proves to be higher than carefully chosen baselines, and it works best for heterogeneous graphs. Future works could examine special cases of Greedy DMC, as pointed out before. We also provided a brief overview of synthetic graph generation models and a short discussion of the difficulty of modeling complex networks.
7 Acknowledgements
We thank Allison Gunby-Mann for introducing Ashley to the graph alignment problem and bringing the random walk sampling method to attention.
Appendix A Examples to Demonstrate DMC
A.1 Degree Matrix Formulation
Let us examine the graph (assumed to be a local picture taken from a much larger graph) in Figure 8 with circled node as the origin. Then the boxed nodes are the first neighbors of the origin. Starting from the first neighbor in the upper left corner and going counter clock-wise, the degrees of the five first neighbors are $3,4,3,5,3$. Suppose (arbitrarily) $m$ in this case is $10$, then the row vector representation for the circled node would be $[3,3,3,4,5,0,0,0,0,0]$ in the degree matrix for the larger graph. We do this for every node in the larger graph which this smaller piece is taken from to build a degree matrix.
Figure 8: Example Graph to Demonstrate Degree Matrix Formulation
A.2 DMC Demonstration
Now consider the two graphs in Figure 9 as a pair that we are to align as a simple example.
(a) $F_{1}$
(b) $F_{2}$
Figure 9: Example Graphs to Demonstrate DMC
Let the graph in Figure 9 (a) be $F_{1}=(V_{F_{1}},E_{F_{1}})$ and the graph in Figure 9 (b) be $F_{2}=(V_{F_{2}},E_{F_{2}})$. The node set $V_{l}$ of the largest overlapping subgraph consists of all five nodes, or in mathematical notation: $V_{l}=V_{F_{1}}=V_{F_{2}}$. However, the graphs are not isomorphic, so the edge set $E_{l}$ of the overlapping subgraph is a strict subset of $E_{F_{1}},E_{F_{2}}$, which in mathematical notation is: $E_{l}\subsetneq E_{F_{1}},E_{F_{2}}$. Let the edge between node 3 and node 4 in $F_{2}$ be $w$. Then the overlapping subgraph takes the shape of $F_{2}\backslash\{w\}$. The degree matrices for $F_{1}$ and $F_{2}$ are
[…]
Then we use the Munkres assignment algorithm to minimize the cost of row alignment between the two matrices above. The results indicate that the algorithm aligns nodes $1,2,3,4,5$ of $F_{1}$ to nodes $2,5,1,3,4$ of $F_{2}$ (one could check understanding of the concepts with this result). This is not an optimal result, but it is expected since the graphs in this example have low heterogeneity. This example is purely for the purpose of demonstrating the method.
Appendix B Munkres Assignment Algorithm
Algorithm 1 Munkres Assignment Algorithm
1:  Input: A cost matrix $C$ of size $n\times n$.
2:  Output: An optimal assignment and the total cost.
3:  Step 1: Subtract the row minimum.
4:  for each row $i$ in $C$ do
5:     Subtract the minimum value of row $i$ from all elements in row $i$.
6:  end for
7:  Step 2: Subtract the column minimum.
8:  for each column $j$ in $C$ do
9:     Subtract the minimum value of column $j$ from all elements in column $j$.
10:  end for
11:  Step 3: Cover all zeros with a minimum number of lines.
12:  while not all zeros are covered do
13:     Identify rows and columns containing uncovered zeros.
14:     Cover all zeros using the minimum number of horizontal and vertical lines.
15:  end while
16:  Step 4: Check the number of lines.
17:  if the number of lines is equal to $n$ then
18:     An optimal assignment is possible.
19:     Go to Step 6.
20:  else
21:     Go to Step 5.
22:  end if
23:  Step 5: Adjust the matrix.
24:  Find the smallest uncovered value.
25:  Subtract this value from all uncovered elements.
26:  Add this value to elements at intersections of covering lines.
27:  Return to Step 3.
28:  Step 6: Make the assignments.
29:  Select zeros such that no two are in the same row or column.
30:  Return the assignment and compute the total cost.
Appendix C Auction Algorithm
Algorithm 2 Auction Algorithm with Munkres for Unassigned Items
0:  Profit matrix $P$ of size $n\times m$ (agents $\times$ tasks), threshold $\epsilon>0$.
0:  Assignment of agents to tasks.
1:  Initialize $prices[j]\leftarrow 0,\;\forall j\in\{1,\ldots,m\}$.
2:  Initialize $assignments[i]\leftarrow\text{unassigned},\;\forall i\in\{1,\ldots,n\}$.
3:  while any agent $i$ is unassigned do
4:     for each unassigned agent $i$ do
5:        Compute net values $v[j]\leftarrow P[i,j]-prices[j],\;\forall j$.
6:        Identify the best task: $j_{1}\leftarrow\arg\max_{j}v[j]$.
7:        Identify the second-best task: $j_{2}\leftarrow\arg\max_{j\neq j_{1}}v[j]$.
8:        Compute bid increment: $\Delta\leftarrow v[j_{1}]-v[j_{2}]+\epsilon$.
9:        Update the price: $prices[j_{1}]\leftarrow prices[j_{1}]+\Delta$.
10:        Assign task $j_{1}$ to agent $i$.
11:        if task $j_{1}$ was previously assigned to agent $k$ then
12:           Unassign agent $k$.
13:        end if
14:     end for
15:  end while
16:  if any tasks remain unassigned then
17:     Assign remaining tasks using the Munkres assignment algorithm.
18:  end if
19:  Return: Final assignment of agents to tasks.

### v3

@@ 5.1.1 Real World Networks
In Table 1, we display statistics that compare DMC and Greedy DMC with baselines after applying all methods to the unweighted multi-layer PPI network [10, 12, 23, 31]. Using the edge deletion sampling method, we set deletion probability $p_{d}=0.01$. Regarding the different graph alignment methods, we note that the Greedy DMC uses a threshold of $\epsilon=100$. “Two-Step” is, as its name suggests, a two-step correction method that uses degrees for alignment, and refines that first-step alignment through average degrees of neighbors. “Euc Dist.” aligns two graphs by minimizing Euclidean distances between embeddings. We can observe through direct comparison that DMC has highest accuracy, but larger complexity compared to REGAL and FINAL [12, 31]. The Greedy DMC strikes a similar balance compared to FINAL: while it is more accurate, it could also be higher in time complexity.
Table 1: Comparison with baselines on PPI network.
Method Score (%) Complexity
DMC 96.09 $\sim O(n^{2})-O(n^{3})$
Greedy DMC 50.46 $\sim O(n^{2})-O(n^{3})$
REGAL $9080<p<$ 90 $\sim O(n)-O(n^{2})$
FINAL $4535<p<$ 45 $\sim O(n^{2})$
Klau $3020<p<$ 30 $\sim O(n^{5})$
IsoRank $155<p<$ 15 $\sim O(n^{4})$
Two-Step 4.42 $\sim O(n^{2})-O(n^{3})$
Euc Dist. 0.15 (near zero) $\sim O(n^{2})-O(n^{3})$
After showing the potential of DMC and Greedy DMC by comparing to baselines on a classic network, we turn to demonstrating that DMC works better when networks are more heterogeneous. We first draw an observed correlation between variance in degree distribution and performance of the DMC; the correlation is not strict, but implies from at least one perspective that the DMC works well for heterogeneous networks in general. Table 2 provides a summary of statistics for social networks and roadmaps, and Figures 6 and 7 provides results applying DMC to the networks. The best performing Facebook and YouTube have large variances in their degree distributions. Comparing Flickr, Last.fm, and Myspace, Flickr and Last.fm are more suitable for DMC, which is most likely due to their much larger variance in degree distribution. However, the correlation between variance and performance can be false at times, as can be seen from the California and Pennsylvania roadmaps. The two roadmaps have close degree distribution means and variances, with the Pennsylvania one having lower degree variance, but it outperforms the California roadmap significantly.
Table 2: Means and variances of degree distributions of $G_{s}$.
@@
Figure 6: Social Networks
Figure 7: Roadmaps
Moreover, we show that DMC works better on graphs with high densities and clusterings in Table 3. We provide results of applying DMC to more biological networks randomly chosen from the Brain Network and Biological Network sections in the Network Repository database. In Table 3, “grid-fission-yeast”, “ce-gn”,“grid-human”, and “yeast” are from the Biological category, and “mouse-retina-1”, “fly-drosophila-medulla-1”, and “mouse-kasthuri” are from the Brain category.
Table 3: Table of experiments on biological networks, ordered by DMC’s performance on $p=90\%$ overlap subgraphs.
Networks Score Density Clustering
grid-fission-yeast .9932 .0123 .1874
mouse-retina-1 .9900 .9983 1.5539
ce-gn .9894 .0218 .1839
fly-drosophila-medulla-1 .9378 .0211 3.8729
grid-human .8440 .0014 .0800
mouse-kasthuri .6156 .0032 0
yeast .5822 .0018 .0708
We note that the highlighted densities are those above the 0.01 threshold, and for the average clustering coefficient the threshold is 0.1 (density and clustering taken from Network Repository). The graphs with both a high density and a high clustering coefficient outperformed the other graphs significantly.
In this paragraph, we examine the behavior of DMC and Greedy DMC applied to the unweighted multi-layer PPI network under different conditions. For DMC, we see in Figure 8 that DMC’s performance decays as deletion probability (noise level) increases from 0.01 to 0.1. This is expected and no surprise for a new algorithm, but it is worth pointing out that the performance does not rapidly drop to near zero values even with ten times the experimental noise level, implying that the method has potential to withstand noise. In Table 4, we see that when $\epsilon=0,10$, the Greedy DMC algorithm maintains high performance. We will study these special cases in the future.
Figure 8: Performance Decay Plot
Table 4: Greedy DMC Performance
$\epsilon$ Score (%)
0 96.58%
10 96.32%
100 50.46%
1000 0.18%
5.1.2 Artificial Networks
In this section, we introduce the Erdös-Rényi, Barabási-Albert, and Chung-Lu models that were derived to generate complex networks, potentially to model real world network behaviors.
The Erdös-Rényi model is the most basic random graph generation model. One common way to define this is to fix the number of vertices, say $n_{er}$, then connect an edge between every pair of vertices with some probability $p_{er}$. The resulting graph is typically denoted as $G(n_{er},p_{er})$ [7]. One interesting property of these graphs is that their degree distributions are binomial, as can be be seen by the following calculation:
@@ 5.1.2 Artificial Networks
The Barabási-Albert model considers properties observed on real world networks, especially the preferential attachment phenomena [2]. Instead of attaching new edges with uniform probability, it is possible to attach new edges with weighted probabilities. Suppose we already have $m$ nodes with some connections, then for the $(m+1)^{\text{th}}$ node, the probability that it will be attached to a previous node $v_{i}$ is
@@
We can observe that the probabilities of attaching the new node to any previous node sums up to one, forming a probability distribution; moreover, higher degree nodes attract more new nodes. This usually leads to what we call hubs, which are local centers that are attached to many low degree nodes. It is common for hubs to exist in biological and social networks, and hubs add to the heterogeneity of graphs. However, having hubs with too much concentration could lead to confusion when differentiating low-degree nodes. Therefore, it is always good to strike a balance between concentrating and diffusing nodes, but how this is done in the real world is yet unknown.
Now we introduce the Chung-Lu model. It also attempts to capture real world network properties, but instead of focusing on preferential attachment or community behaviors, they attempt to directly approximate real world networks’ degree distributions [5]. Suppose we have a specific type of degree distribution in mind and we expect $m$ nodes in total, then the attachment probability between each node $v_{i}$ and another node $v_{j}$ is as follows:
@@
In (8), the degrees are target degrees according to our degree distribution in mind. With the attachment process described above, the expected degree at each node $v_{i}$ is
@@
Despite the careful construction of models, human knowledge of real world networks are nowhere near complete. Through comparison of DMC results on real world networks with experimental results on synthetic graphs, we will see that the complexity and irregularity of real world networks make their nodes more effectively differentiable through the DMC, and there is still a notable gap between the synthetic models and real world networks since they exhibit different behaviors in response to the same graph alignment method.
@@
Definition 5 (k-partite Graph).
@@
We first applied DMC to graphs generated through the Erdös-Rényi and Barabási-Albert models. We examine the performance of DMC on the graphs with respect to different partite numbers $k$. For our Erdös-Rényi experiments, we fix total number of nodes to be $n_{er}=100$, then split the nodes into $k$ partite groups; the partition of nodes is arbitrary. After that, we connect edges between different partite groups with probability of $p_{er}=0.5$. Then we use the random walk graph sampling method to obtain two graphs with 90% overlap, and perform DMC. We run the same experiment for 20 times, then take the mean performance. For our Barabási-Albert experiments, we do a similar thing. Every time we add a new node, we connect it to five different previous nodes (add 5 edges).
We not only examined the completely random Erdös-Rényi and heterogeneous Barabási-Albert, but also experimented with intermediate graphs (i.e. a spectrum of graphs ranging from most uniform to most heterogeneous). We move along the spectrum of Poisson distributions (see Figure 11) with different parameter values, which can approximate distributions from binomial (black) to power law (yellow). A formal representation of the Poisson distribution for each node $v$ is as follows:
[…]
where $\lambda$ is a parameter. These pre-determined distributions are encoded in graphs through the Chung-Lu model. Figures 12, 13, 14, and 15 show how the behavior of DMC converges to the heterogeneous end.
@@
Real world networks are more complex than we can imagine, hence the difficulty of capturing them with mathematical language. However, it is interesting the simple DMC works better on the more complex real world networks than synthetically generated graphs. This shows that modeling complex networks is no easy task, and understanding the complexity will potentially bring us rich information. Moreover, the fact that DMC has higher performance when using the more sophisticated Barabási-Albert and Chung-Lu, instead of Erdös-Rényi, indicates that it is no coincidence that DMC works well on heterogeneous real world networks. Additionally, while the Erdös-Rényi model was not as suitable for DMC, we discovered as a side result a linear relationship between the performance of DMC and the partite number of the Erdös-Rényi networks.
Figure 9: Erdos-Renyi Model: negative linear relationship.
Figure 10: Barabasi-Albert Model
Figure 11: Evolution of Poisson distribution with decreasing $\lambda$. Taken from Scribbr.
Figure 12: Performance of DMC applied to Chung-Lu generated graphs with respect to different partite numbers when parameter $\lambda=20$.
Figure 13: Performance of DMC applied to Chung-Lu generated graphs with respect to different partite numbers when parameter $\lambda=10$.
Figure 14: Performance of DMC applied to Chung-Lu generated graphs with respect to different partite numbers when parameter $\lambda=5$.
Figure 15: Performance of DMC applied to Chung-Lu generated graphs with respect to different partite numbers when parameter $\lambda=1$.
5.2 Weighted Graphs: Brief Examination of Weighted DMC
As previously mentioned, we also dedicate a brief section to the Weighted DMC to demonstrate its potential, which in turn will also support that DMC, albeit its simplicity, is a valid approach. Table 5 provides the results of applying the Weighted DMC to five weighted PPI networks for different species that are typical subjects of biological studies, all taken from the STRING dataset [26]. Among the five species, E. Coli is the only prokaryotic organism, meaning that it has simpler cell structure lacking a sophisticated nucleus. We can reasonably speculate that the PPI for E. Coli is less heterogeneous, leading to a less accurate Weighted DMC result. We have also discussed in Section 5.1 that heterogeneous networks tend to be more suitable for the simple DMC. The other four PPI networks all yielded performance above $95\%$, which is high.
We run the entire process, from random walk to graph alignment, for ten times and take their mean performance. All experiments sampled 5000-node graphs except for E. Coli, where we sampled 3000-node graphs, due to its limited size. We used $p_{d}=0.01$ deletion probability for all experiments.
Table 5: Weighted PPI Networks
Species Score (%)
Human 99.19
Yeast 98.06
Mouse 97.27
Fruit Fly 96.26
E. Coli 91.46
We have also run experiments on social networks, where the results are reported in Table 6. “soc-sign-bitcoin-otc” and “soc-sign-bitcoin-alpha” are taken from the SNAP database [17], and “music-cotagging”, “ai-cotagging”, “apple-cotagging”, “travel-cotagging”, and “economics-cotagging” are taken from the Computer Science Department webpage of Cornell University [9]. The “cotagging” networks are networks by keywords on Stack Exchange. We used the edge deletion sampling method, setting deletion probability $p_{d}=0.01$, and run all experiments ten times to take the mean performance. We can speculate here that including weight information of edges has helped us gain more insight into local graph structure, since the overall performance of the Weighted DMC on the social networks is higher than that of DMC applied to unweighted social networks.
Table 6: More Weighted Networks
Network Score (%) Sampled Nodes
soc-sign-bitcoin-otc 77.44 5000
soc-sign-bitcoin-alpha 82.08 3000
music-cotagging 98.91 467
ai-cotagging 96.95 285
apple-cotagging 98.65 1063
travel-cotagging 97.35 1779
economics-cotagging 98.27 369
@@ 6 Conclusion and Future Work
We proposed two graph alignment methods for unweighted networks: DMC and Greedy DMC, and also another one for weighted networks, which we named Weighted DMC. We showed that they were theoretically motivated, and analyzed experimental results. DMC’s accuracy on classic PPI networks proves to be higher than carefully chosen baselines, and it works best for heterogeneous graphs which has been demonstrated from multiple perspectives. This process included providing an overview of synthetic graph generation models and a short discussion of the difficulty of modeling complex networks. Future works could examine special cases of Greedy DMC and the Weighted DMC, as pointed out in previous sections; one could also test this sequence of methods with more datasets to more clearly define the optimal set of networks that should be aligned using DMC, a simple method that can easily be replicated.